OpenStudy (anonymous):

evaluate: log[6](216) a. 1/3 b.4 c.3 d.1/6 e. none

OpenStudy (anonymous):

C. \[6^3=216\]or\[\frac{\ln(216)}{\ln(6)}=3\]

OpenStudy (anonymous):

thank you